OpenStudy (anonymous):

What is the slope of the line that passes through the points (3, –1) and (–2, –5)? 1. -4 2. -1/4 3. 1/4 4. 4

OpenStudy (anonymous):

Huh, I got 4/5. Maybe I'm just that off.

OpenStudy (anonymous):

m = slope = (y2-y1)/(x2-x1) = (-5 - -1)/(-2 -3) = -4/(-5) ? = 4/5 ME TOO

OpenStudy (anonymous):

OMG sorry this s the question i gave you guys the wrong one my bad :3 What is the slope of the line passing through points A and B? http://postimg.org/image/6q19uycmb/

OpenStudy (anonymous):

1/4

OpenStudy (anonymous):

Thanks OrangeMaster :3